# ltxutilRepair the float

### v1

@@ Appendix B Assignment strategies for surface codes
strategy iii: random Errors with strong correlation are similar to burst errors in classical communication in the sense that the errors have spatial locality. This locality of errors can be addressed by classical error-correcting codes using two methods. The first method treats multiple bits as a single symbol (an element of a finite field), such as BCH codes and Reed-Solomon codes [62]. Thanks to the high ability to correct burst errors, Reed-Solomon codes are used in many classical systems, including QR codes [63], CDs, and satellite communications. Another method is the interleaving [64] technique. Interleaving eliminates locality by permuting the rows and columns of the code’s generator matrix. There is also a method to apply interleaving to QECCs[65]. Inspired by interleaving, we have constructed two strategies for quantum multiplexing with randomness.
strategy iv: random + threshold The fourth strategy is a modified version of the third strategy. The pseudo-code is shown below in Algorithm 1. The flow of the algorithm is as follows: A “threshold” $T$ is set as $12/d-$ 1, which is the maximal distance between two qubits in the $[\![2d^{2},2,d]\!]$ toric codes. This value will be used to check that the set of qubits in the same photon has enough distance between each other. Then, it randomly assigns qubits for each photon while respecting the distance threshold. It randomly selects the first qubit of the photon, then it randomly selects a qubit again and takes it as a candidate to assign it to this photon. When the distance between the candidate qubit and the qubit(s) already in the photon is greater than the threshold, the qubit is accepted, and when it is less, it is rejected. This procedure is repeated until the photon has been fulfilled. If no qubit satisfies the threshold, the threshold value is lowered by one. This is to ensure that the algorithm will always finish running. By repeating this process, we can assign all the qubits to photons. This strategy is designed to have randomness and to increase the distance between qubits in the same photon.
This algorithm requires calculating the distance between two qubits, which is easy for the surface codes because the taxicab metric defines the distance (Manhattan distance). Note that this and other assignment strategies can still be applied even if the number $m$ of qubits per photon is not a divisor of the total number of qubits. In this case, we allow for a final “remainder” photon containing fewer than $m$ qubits.
Input: $P=\{p_{i}\}$ (the set of photons) where initially $p_{i}=\{\emptyset\}$ (the set of qubits to be encoded in the $i^{\text{th}}$ photon), $Q=\{q_{j}\}$ (the list of all physical qubits in the code), and the number $m$ of qubits in a single photon.
Output: $P=\{p_{i}\}$ (set of set of qubits in $i^{\text{th}}$ photon).
@@
4       Pick a qubit $q_{j}\in Q$ randomly.;
@@ 5       Move $q_{j}$ from $Q$ to $p_{i}$;
8                   Pick a candidate qubit $q_{k}\in Q$ randomly;
9                   if $q_{k}$ has minimum distance greater than $T$ from all the qubits in $p_{i}$ then
@@ 16             Update $T$ := $T-1$;
Appendix C Peeling Decoder
@@
C.1 Peeling Algorithm for Classical Codes
@@ 17             Return Failure;
C.2 Peeling Algorithm for CSS Codes
@@
C.3 Peeling Algorithm for Surface Codes
@@
This modified peeling algorithm is a linear-complexity, maximum-likelihood decoder for the surface code. We use this algorithm in our numerical simulations for the surface code. Our implementation of the surface code peeling decoder is available in [57]. This process is briefly summarized in Fig. 16 and 17.
Figure 16: Illustration of the error correction process for an example of erasure errors with the surface code lattice. (1) Erased qubits are shown in bold grey lines. (2) Erasure errors are converted to random Pauli errors by replacing erased qubits with mixed states. The syndrome (indicated by the red vertices) is then computed by applying stabilizer measurement as explained in Sec. 2.2. (3) Information seen by the decoding algorithm: erasure pattern and syndrome. (4) A spanning tree for the erasure pattern in the lattice is computed; this is identified with a corresponding acyclic subgraph of the Tanner graph. The surface code peeling decoder then corrects the qubits one by one using this subgraph.
@@
C.4 Peeling Algorithm for HGP Codes
The pruned peeling + VH decoder [59] is yet another generalization of the peeling decoder adapted to the special case of HGP codes. As mentioned in Appendix C.2, because these are a type of CSS code, the standard peeling algorithm can be directly applied to HGP codes. However, this algorithm performs very poorly in practice, even for LDPC codes. This poor performance can be explained by the presence of stopping sets unique to HGP codes which have no analogue in the classical case. These stopping sets can be grouped into two types: stabilizer and classical, both of which cause the decoder to fail. The pruned peeling + VH decoder is designed to address these stopping sets.
Stabilizer stopping sets occur when the erasure contains the qubit-support of an $X$- or $Z$-type stabilizer. Recall that, for a CSS code, $X$- and $Z$-stabilizers overlap on an even number of qubits. Hence, restricting to the Tanner graph of $H_{Z}$, the check-nodes in the subgraph corresponding to the qubit-support of an $X$-stabilizer all have even degrees. In particular, there exist no dangling checks (which have degree 1) and hence this is a peeling decoder-stopping set. A similar relationship holds for $Z$-stabilizers in the Tanner graph of $H_{X}$.
@@ C.4 Peeling Algorithm for HGP Codes
The pruned peeling + VH decoder refers to the combination of these three strategies (standard peeling, correction of stabilizer stopping sets, and correction of classical stopping sets). For simplicity, we also use the term combined decoder to refer to peeling + pruned peeling + VH. A stopping set for the combined decoder meets three conditions: there exist no remaining dangling checks; the remaining erasure does not cover the qubit-support of a stabilizer; remaining classical stopping sets form a cycle. Although these happen infrequently at a low erasure rate, an erasure pattern of this form will result in a decoder failure. These are distinct from logical errors, which can only be identified in numerical simulations where the decoding algorithm successfully terminates. The maximum-likelihood decoder always terminates, and thus, logical errors are the only source of failures. Although we make a distinction between these two possibilities, we will use the term error recovery failure to refer to either a decoder failure or a non-decoder failure logical error. A more detailed discussion of these differences is included in Appendix D.2.
@@
Appendix D Additional Details for HGP Codes
D.1 Symmetric Constructions
@@
D.2 Types of Error Recovery Failures for the Pruned Peeling + VH Decoder
In Sec. 5.2, we observed that the combined decoder (peeling + pruned peeling + VH) is not maximum-likelihood since decoder failures are possible in addtion to logical errors. Failure rate in the literature usually refers to the logical error rate, which is the only source of errors for a maximum-likelihood decoder. Logical errors for the erasure channel can only occur when the erasure covers a logical code word. However, there may exist erasure patterns covering a logical error which result in a decoder failure, and hence are not properly identified as logical errors. This distinction is stated visually by the Venn diagram of Fig. 18. The failure rate computed in our numerical simulations for the combined decoder is the cummulative effect of these two possibilities, what we refer to as error recovery failure rate on the vertical axis in the plots of our numerical simulations for HGP codes. Note that failures at low erasure rates are almost exclusively due to logical errors, and so this distinction can be regarded as negligible in the practical regime.
@@ D.2 Types of Error Recovery Failures for the Pruned Peeling + VH Decoder
Figure 18: Venn diagram distinguishing between the types of failures possible using the pruned peeling + VH decoder. A decoding failure (DF) occurs when the decoder becomes stuck in a stopping set it cannot correct. A logical error (LE) occurs when the decoder successfully terminates with a predicted error, but the actual and predicted errors combine to give a logical code word. An error recovery failure (ERF) refers to the union of these two possibilities.
D.3 Technical Details for Assignment Strategies used with Multiplexed HGP Codes
@@
3       Pick a qubit $q_{j}\in Q$ randomly;
@@ 13      Move all qubits in $Q^{\prime}$ back to $Q$;
15             Pick a qubit $q_{k}\in Q$ randomly;

### v2

@@ Appendix C Assignment strategies for surface codes
strategy iii: random Errors with strong correlation are similar to burst errors in classical communication in the sense that the errors have spatial locality. This locality of errors can be addressed by classical error-correcting codes using two methods. The first method treats multiple bits as a single symbol (an element of a finite field), such as BCH codes and Reed-Solomon codes [63]. Thanks to the high correctability of burst errors, Reed-Solomon codes are used in many classical systems, including QR codes [64], CDs, and satellite communications. Another method is the interleaving [65] technique. Interleaving eliminates locality by permuting the rows and columns of the code’s generator matrix. There is also a method to apply interleaving to QECCs[66]. Inspired by interleaving, we have constructed two strategies for quantum multiplexing with randomness.
strategy iv: random + threshold The fourth strategy is a modified version of the third strategy. The pseudo-code is shown below in Algorithm 1. The flow of the algorithm is as follows: A “threshold” $T$ is set as $12/d-$ 1, which is the maximum distance between two qubits in the $[\![2d^{2},2,d]\!]$ surface code. This value will be used to check that the distance between the qubits in the same photon is high enough. Then, it randomly assigns qubits for each photon while respecting the distance threshold. It randomly selects the first qubit of the photon, then it randomly selects a qubit again and takes it as a candidate to assign it to this photon. When the distance between the candidate qubit and the qubit(s) already in the photon is greater than the threshold, the qubit is accepted, and when it is less, it is rejected. This procedure is repeated until the photon has been assigned the appropriate number of qubits. If no qubit satisfies the threshold, then the threshold value is lowered by one. This ensures that the algorithm will always terminate. By repeating this process, we can assign all the qubits to photons. This strategy is designed to leverage randomness while also increasing the distance between qubits in the same photon.
This algorithm requires calculating the distance between two qubits, which is easy for the surface code thanks to the fact that the distance between qubits in the grid representation is defined by the taxicab metric (the Manhattan distance). Note that this and other assignment strategies can still be applied even if the number $m$ of qubits per photon is not a divisor of the total number of qubits. In this case, we allow for a final “remainder” photon containing fewer than $m$ qubits.
Input: $P=\{p_{i}\}$ (the set of photons) where initially $p_{i}=\{\emptyset\}$ (the set of qubits to be encoded in the $i^{\text{th}}$ photon), $Q=\{q_{j}\}$ (the list of all physical qubits in the code), and the multiplexing number $m$ of qubits in each photon.
Output: $P=\{p_{i}\}$ (a set of sets of qubits assigned to the $i^{\text{th}}$ photon).
@@
4       Pick a qubit $q_{j}\in Q$ randomly;
@@ 5       Move $q_{j}$ from $Q$ to $p_{i}$;
8                   Pick a candidate qubit $q_{k}\in Q$ at random;
9                   if the distance between $q_{k}$ and all qubits in $p_{i}$ is greater than $T$ then
@@ 16             Update $T$ := $T-1$;
Appendix D Peeling Decoder
@@
D.1 Peeling Algorithm for Classical Codes
@@ 17             Return Failure;
D.2 Peeling Algorithm for CSS Codes
@@
D.3 Peeling Algorithm for Surface Codes
@@
This modified peeling algorithm is a linear-complexity, maximum-likelihood decoder for the surface code. We use this algorithm in our numerical simulations for the surface code. Our implementation of the surface code peeling decoder is available in [57]. This process is briefly summarized in Fig. 17 and 18.
Figure 17: Illustration of the error correction process for an example of erasure errors with the surface code lattice. (1) Erased qubits are shown in bold grey lines. (2) Erasure errors are converted to random Pauli errors by replacing erased qubits with mixed states. The syndrome (indicated by the red vertices) is then computed by applying stabilizer measurements as explained in Sec. 2.2. (3) Information seen by the decoding algorithm: erasure pattern and syndrome. (4) A spanning tree for the erasure pattern in the lattice is computed; this is identified with a corresponding acyclic subgraph of the Tanner graph. The surface code peeling decoder then corrects the qubits one by one using this subgraph.
@@
D.4 Peeling Algorithm for HGP Codes
The pruned peeling + VH decoder [60] is yet another generalization of the peeling decoder adapted to the special case of HGP codes. As mentioned in Appendix D.2, because these are a type of CSS code, the standard peeling algorithm can be directly applied to HGP codes. However, this algorithm performs very poorly in practice, even for LDPC codes. This poor performance can be explained by the presence of stopping sets unique to HGP codes which have no analogue in the classical case. These stopping sets can be grouped into two types: stabilizer and classical, both of which cause the decoder to fail. The pruned peeling + VH decoder is designed to address these stopping sets.
Stabilizer stopping sets occur when the erasure contains the qubit-support of an $X$- or $Z$-type stabilizer. Recall that the $X$- and $Z$-type stabilizers must commute, which is possible if and only if they overlap on an even number of qubits. Hence, in terms of the nodes denoting stabilizer generators in the Tanner graph, if we fix a node corresponding to an $X$-stabilizer generator and consider the set of qubits on which it is supported, each $Z$-stabilizer generator with overlapping qubit-support must overlap on an even number of qubits in order to satisfy the commutativity condition. Since the nodes denoting $Z$ stabilizer generators are exactly the check nodes in the Tanner graph of $H_{Z}$, these constitute a stopping set, as shown in the example of Fig. 19. In particular, there exist no dangling checks (which have degree 1) and hence this is a peeling decoder-stopping set. A similar relationship holds for $Z$-stabilizers in the Tanner graph of $H_{X}$.
$n_{1}$
$n_{2}$
$r_{1}$
$r_{2}$
Figure 19: An example of a stabilizer stopping set for $H_{z}$ of the $2\times 2$ surface code. The red circles indicate an erasure error, which consists of all of the qubits adjacent to the $X$-stabilizer denoted by the orange square. The cyan squares denote the $Z$-stabilizers adjacent to this set of erased qubits. Notice that each of these $Z$-stabilizers is adjacent to an even number of erased qubits, which is a consequence of the commutation relation between $X$- and $Z$-type stabilizers and also guarantees that there are no peelable degree-$1$ checks in the erasure subgraph.
@@ D.4 Peeling Algorithm for HGP Codes
The pruned peeling + VH decoder refers to the combination of these three strategies (standard peeling, correction of stabilizer stopping sets, and correction of classical stopping sets). For simplicity, we also use the term combined decoder to refer to peeling + pruned peeling + VH. A stopping set for the combined decoder meets three conditions: there exist no remaining dangling checks; the remaining erasure does not cover the qubit-support of a stabilizer; remaining classical stopping sets form a cycle. Although these happen infrequently at a low erasure rate, an erasure pattern of this form will result in a decoder failure. These are distinct from logical errors, which can only be identified in numerical simulations where the decoding algorithm successfully terminates. The maximum-likelihood decoder always terminates, and thus, logical errors are the only source of failures. Although we make a distinction between these two possibilities, we will use the term error recovery failure to refer to either a decoder failure or a non-decoder-failure logical error. A more detailed discussion of these differences is included in Appendix E.2.
@@
Appendix E Additional Details for HGP Codes
E.1 Symmetric Constructions
@@
E.2 Types of Error Recovery Failures for the Pruned Peeling + VH Decoder
In Sec. 5.2 and Sec. D.4, we observed that the combined decoder (peeling + pruned peeling + VH) is not maximum-likelihood since decoder failures are possible in addtion to logical errors. Failure rate in the literature usually refers to the logical error rate, which is the only source of errors for a maximum-likelihood decoder. Logical errors for the erasure channel can only occur when the erasure covers a logical code word. However, there may exist erasure patterns covering a logical error which result in a decoder failure, and hence are not properly identified as logical errors. This distinction is stated visually by the Venn diagram of Fig. 20. The failure rate computed in our numerical simulations for the combined decoder is the cummulative effect of these two possibilities, what we refer to as error recovery failure rate on the vertical axis in the plots of our numerical simulations for HGP codes. Note that failures at low erasure rates are almost exclusively due to logical errors, and so this distinction can be regarded as negligible in the practical regime.
@@ E.2 Types of Error Recovery Failures for the Pruned Peeling + VH Decoder
[…]
E.3 Technical Details for Assignment Strategies used with Multiplexed HGP Codes
@@
3       Pick a qubit $q_{j}\in Q$ at random;
@@ 13      Move all qubits in $Q^{\prime}$ back to $Q$;
15             Pick a qubit $q_{k}\in Q$ at random;